# Provably Efficient Multi-Objective Bandit Algorithms under Preference-Centric Customization

### v1

@@ 6 The Case with Hidden Preference
Figure 3: A scenario of user’s preference feedback is not explicitly provided.
@@
Next, we consider another practical scenario where only feedback on the reward and overall reward is observable, while preference feedback is not provided. For instance, in hotel surveys, customers often provide ratings on specific objectives (e.g., price, location, environment, amenities) along with an overall rating (as depicted in Fig. 3). In such cases, user preferences can be inferred from the latent relationship between the overall rating and the individual objective ratings. Formally, in each round $t$, the learner selects an arm $a_{t}\!\in\![K]$, and observes the reward vector $\boldsymbol{r}_{a_{t}}\!\in\!\mathbb{R}^{D}$, and the corresponding overall-reward score:
$g_{a_{t},t}=\Phi(\boldsymbol{c}_{t},\boldsymbol{r}_{a_{t},t})=\boldsymbol{c}_{t}^{\top}\boldsymbol{r}_{a_{t},t}\in\mathbb{R}$ (5)
@@
Within this framework, we adhere to the original Assumption 3.1 regarding rewards. It is worth noting that, in many real-world applications like hotel rating systems, the overall rating often shares the same scale as individual objective ratings. Therefore, we assume in this problem that the bound on the overall reward is identical to that of the individual rewards. This introduces one additional assumption and one revised assumption, as outlined below:
Assumption 6.1.
For $t\in[T]$, $a_{t}\in[K]$, the overall-reward score satisfies $g_{a_{t},t}\in[0,1]$.
Assumption 6.2.
@@
For $t\in[T]$, the stochastic preference is bounded and satisfies $\|\boldsymbol{c}_{t}\|_{1}\leq 1$. Without loss of generality, we assume $\boldsymbol{c}_{t}(d)$ is $R$-sub-Gaussian [^1] By Hoeffding’s lemma, for any $X\in[a,b]$ almost surely, $X$ is a $R$-sub-Gaussian random variable with $R$ at most $(b-a)/2$., $\forall d\in[D]$.
@@ 6.1 Unique Challenges
This problem introduces two unique challenges that distinguish us from previous MAB studies:
Local Exploration vs Global Exploration. Unlike traditional bandit algorithms that focus on a single goal (e.g., identifying the arm with the highest reward), the uncertainty in both preference and reward, combined with the need to infer latent preference, introduces a novel trade-off challenge: balancing global exploration for better preference estimation and local exploration of arm rewards:
@@
Note that these two learning objectives may conflict, as arms with high rewards might lack sufficient information for latent preference learning and could even degrade estimation performance (further discussed in the second challenge). This can also be verified by Fig. 4, where 80 samples of $[\boldsymbol{r}_{t},g_{t}]$ are collected by repeatedly pulling an arm, and preference $\boldsymbol{\hat{c}}$ is estimated using linear regression. Here, $\boldsymbol{c}_{t}$ at each step follows a Gaussian distribution with a mean of $[0.5,0.5]$. The results demonstrate that samples from suboptimal Arm-1 (Fig. 4a) significantly outperform those from Pareto-optimal Arm-2 (Fig. 4b) in preference estimation. This necessitates an exploration policy that effectively addresses both global and local learning objectives.
Random Mapping from $\boldsymbol{r}_{t}$ to $g_{t}$. In the hidden preference case, the observed overall rewards are generated through a random mapping of rewards. Specifically, $g_{a_{t},t}=(\boldsymbol{\overline{c}}+\boldsymbol{\zeta}_{t})^{\top}\boldsymbol{r}_{a_{t},t}=\boldsymbol{\overline{c}}^{\top}\boldsymbol{r}_{a_{t},t}+\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$, where $\boldsymbol{\zeta}_{t}=\boldsymbol{c}_{t}-\boldsymbol{\overline{c}}\in\mathbb{R}^{D}$ is an independent random noise vector. This formulation implies that the overall residual noise term $\zeta_{g,t}=\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$ is no longer independent of the input. Consequently, standard regression models become infeasible for preference estimation, as they rely on the assumption that the residual noise in the output is independent of the input.
Additionally, the magnitude of overall residual noise is a monotonically non-decreasing function w.r.t each reward objective, i.e., $\|\zeta_{g,t}(\boldsymbol{r}_{i})\|\leq\|\zeta_{g,t}(\boldsymbol{r}_{j})\|$ iff $\boldsymbol{r}_{i}\preceq\boldsymbol{r}_{j}$. This explains why suboptimal arms typically outperform Pareto-optimal arms for preference estimation in Fig. 4, as selecting profitable arms tends to amplify the residual error, thereby degrading preference learning. Thus, a tailored latent preference estimator is essential to mitigate the expanding error w.r.t the reward and ensure effective preference learning.
@@ 6.2 Our Algorithm
To this end, we propose a novel PRUCB-HP method (Algorithm 2) involving two key designs as follows.
Key design I: WLS-Preference Estimator. As we have seen before, the randomness of preference $\boldsymbol{c}_{t}$ leads to the overall residual noise $\zeta_{g,t}$ be a function w.r.t, input reward $\boldsymbol{r}_{t}$. Moreover, larger input rewards $\boldsymbol{r}_{t}$ result in greater corruption from the residual noise. To resolve this, we employ a weighted least-squares (WLS) estimator for preference learning. Specifically, our algorithm assigns a weight $w_{t}$ to each observed sample and estimates the unknown preference using weighted ridge regression:
$\hat{\boldsymbol{c}}_{t}\xleftarrow{}\operatorname*{arg\,min}_{\boldsymbol{c}\in\mathbb{R}^{D}}\lambda\|\boldsymbol{c}\|_{2}^{2}+\sum_{\ell=1}^{t-1}w_{\ell}({\boldsymbol{c}}^{\top}\boldsymbol{r}_{a_{\ell},\ell}-g_{a_{\ell},\ell})^{2},$
@@
$\boldsymbol{\hat{c}}_{t}=\boldsymbol{V}_{t-1}^{-1}\sum_{\ell=1}^{t-1}w_{\ell}g_{a_{\ell},\ell}\boldsymbol{r}_{a_{\ell},\ell},$ (6)
where the Gram matrix $\boldsymbol{V}_{t-1}=\lambda\boldsymbol{I}+\sum_{\ell=1}^{t-1}w_{\ell}\boldsymbol{r}_{a_{\ell},\ell}\boldsymbol{r}_{a_{\ell},\ell}^{\top}$.
Algorithm 2 Preference UCB with Hidden Preference (PRUCB-HP)
Parameters: $\alpha$, $\lambda$, $\beta_{t}$, $\omega$.
Initialization: $\boldsymbol{\hat{r}}_{i,1}\!\leftarrow\![0]^{D},N_{i,1}\!\leftarrow\!0,\forall i\!\in\![K]$, $\boldsymbol{\hat{c}}_{1}\!\leftarrow\![\frac{1}{D}]^{D}$, $\boldsymbol{V}_{0}\!\leftarrow\!\lambda\boldsymbol{I}$.
for $t=1,\cdots,T$ do
     Compute reward bonus term: $B_{i,t}^{r}=\|\hat{\boldsymbol{c}}_{t}\|_{1}\sqrt{\frac{\log{t/\alpha}}{\max\{N_{i,t},1\}}},\forall i\in[K]$.
     Compute pseudo information gain term: $B_{i,t}^{c}=\beta_{t}\left\|\hat{\boldsymbol{r}}_{i,t}\!+\!\sqrt{\frac{\log{t/\alpha}}{\max\{N_{i,t},1\}}}\boldsymbol{e}\right\|_{\boldsymbol{V}_{t-1}^{-1}},\forall i\in[K]$.
     Draw arm $a_{t}\leftarrow\operatorname*{arg\,max}_{i\in[K]}\hat{\boldsymbol{c}}_{t}^{\top}\hat{\boldsymbol{r}}_{i,t}+B_{i,t}^{r}+B_{i,t}^{c}$.
     Observe reward $\boldsymbol{r}_{a_{t},t}$ and overall-reward $g_{a_{t},t}$.
     Updating phase:
     $N_{i,t+1}\!=\!N_{i,t}\!+\!\mathds{1}_{\{a_{t}=i\}}$.
     $\hat{\boldsymbol{r}}_{i,t+1}\!=\!(\hat{\boldsymbol{r}}_{i,t}N_{i,t}+\boldsymbol{r}_{a_{t},t}\cdot\mathds{1}_{\{a_{t}=i\}})/N_{i,t+1},\forall i\in[K]$.
     $w_{t}=\frac{\omega}{\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}}$, $\boldsymbol{V}_{t}=\boldsymbol{V}_{t-1}+w_{t}\boldsymbol{r}_{a_{t},t}\boldsymbol{r}_{a_{t},t}^{\top}$.
     $\boldsymbol{\hat{c}}_{t+1}=\boldsymbol{V}_{t}^{-1}\sum_{\ell=1}^{t}w_{\ell}g_{a_{\ell},\ell}\boldsymbol{r}_{a_{\ell},\ell}$.
end for
Inspired by [18] using the inverse of the noise variance as weight for tight variance-dependent regret guarantee, we define the weight as the inverse of the squared $\ell_{2}$-norm of the reward: $w_{t}=\omega/\|\boldsymbol{r}_{a_{t},t}\|_{2}^{2}$, where $\omega>0$ is a threshold parameter guaranteeing $w_{t}\geq 1$. Intuitively, it ensures samples with high rewards will be assigned smaller weights to reduce the influence of potentially large residual noises, while samples with low rewards receive larger weights to ensure their contribution to the estimation.
To see how our choice of weight can tackle the random mapping issue, we first define $\boldsymbol{r}_{a_{t},t}^{\prime}=\sqrt{w_{t}}\boldsymbol{r}_{a_{t},t}$, $g_{a_{t},t}^{\prime}=\sqrt{w_{t}}g_{a_{t},t}$, then the original formula Eq. 5 can be rewrite as
$\sqrt{w_{t}}\cdot g_{a_{t},t}=\sqrt{w_{t}}\cdot\boldsymbol{c}_{t}^{\top}\boldsymbol{r}_{a_{t},t}=\sqrt{w_{t}}\cdot(\boldsymbol{\overline{c}}+\boldsymbol{\zeta_{t}})^{\top}\boldsymbol{r}_{a_{t},t}$
$\implies g_{a_{t},t}^{\prime}=\boldsymbol{\overline{c}}^{\top}\boldsymbol{r}_{a_{t},t}^{\prime}+\sqrt{w_{t}}\cdot\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}.$ (7)
For term $\sqrt{w_{t}}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$, we have the following lemma:
@@
Under Assumption 6.2, the random variable $\sqrt{w_{t}}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t},t}$ is sub-Gaussian with constant $R^{\prime}=\sqrt{\omega}R$.
The proof is available in Appendix D.1. By above lemma, we observe that with the designed weight, the original random mapping regression problem is transferred into a new formula as (7). Specifically, the output $g_{a_{t},t}^{\prime}$ is mapped from $\boldsymbol{r}_{a_{t},t}^{\prime}$ via a fixed vector $\boldsymbol{\overline{c}}$ with a normed $R^{\prime}$-sub-Gaussian residual noise, where $R^{\prime}=\sqrt{\omega}R$, independent of the input.
Key design II: Dual-Exploration Policy. As discussed earlier, there is a new global-local exploration dilemma in our setting. On the one hand, the algorithm must focus on local exploration by selecting optimistically profitable arms to discover better ones. Simultaneously, it must globally explore diverse arms to gather information about the relationship between $\boldsymbol{r}_{t}$ and $g_{t}$ for modeling the hidden preference.
To resolve this, we design an optimistic dual-exploration policy by incorporating a preference-driven bonus and reward-driven bonus within the preference-aware optimization framework for trade-off. The optimistic policy is defined as
$a_{t}\leftarrow\operatorname*{arg\,max}_{i\in[K]}\Phi(\boldsymbol{\hat{c}}_{t},\hat{\boldsymbol{r}}_{i,t})+B_{i,t}^{r}+B_{i,t}^{c},$ (8)
where $\Phi$ is the inner product aggregation function, $B_{i,t}^{r}$ and $B_{i,t}^{c}$ are the dual-exploration bonus terms. We detail the design of these bonus terms below and will later theoretically demonstrate in Section 6.3 how they establish a tight UCB for the expected overall reward, ensuring the effectiveness of the optimistic policy in (8).
Reward Bonus $B_{i,t}^{r}$. The reward bonus term explicitly encourages local exploration of arms with potentially high rewards, for the principle of optimism in face of uncertainty. Specifically, the bonus $B_{i,t}^{r}$ is formulated as a reward uncertainty-aware regularization term:
$B_{i,t}^{r}=\rho_{i,t}^{\alpha}\|\hat{\boldsymbol{c}}_{t}\|_{1}.$ (9)
@@
Preference Bonus $B_{i,t}^{c}$. The preference bonus term aims to encourage the exploration of arms that reduce uncertainty in preference estimation. In previous bandit studies [19, 20, 21] involving linear coefficient ($\theta^{*}$) learning, it has been shown that $\beta\|\boldsymbol{x}_{i}\|_{\boldsymbol{V}^{-1}}$ provides a tight confidence bonus for the payoff of arm $i$, where $\beta$ is the confidence set radius for coefficient estimation, and $\boldsymbol{x}_{i}$ is the observable arm feature. In information theory, $\|\boldsymbol{x}_{i}\|_{\boldsymbol{V}^{-1}}$ also reflects entropy reduction in the model posterior, and is used to measure the estimator uncertainty improvement contributed by the chosen action $\boldsymbol{x_{i}}$ [22].
However, such design is not feasible in our setting, as the exact reward $\boldsymbol{r}_{a_{t},t}$ is revealed only after pulling the arm $a_{t}$, making the actual information gain $\|\boldsymbol{r}_{a_{t},t}\|_{\boldsymbol{V}^{-1}_{t-1}}$ from arm $a_{t}$ unpredictable beforehand. To resolve this problem, we introduce a pseudo information gain term, defined as $\|\boldsymbol{\hat{r}}_{i,t}+\rho_{i,t}^{\alpha}\boldsymbol{e}\|_{\boldsymbol{V}^{-1}_{t-1}}$, where $\rho_{i,t}^{\alpha}$ is the standard Hoeffding bonus. And then the preference bonus is set as
$B_{i,t}^{r}=\beta_{t}\cdot\|\boldsymbol{\hat{r}}_{i,t}+\rho_{i,t}^{\alpha}\boldsymbol{e}\|_{\boldsymbol{V}^{-1}_{t-1}},$ (10)
where the scalar $\beta_{t}$ is the confidence radius of the preference estimation we will give in Lemma 5. Intuitively, this pseudo information gain term captures the potential improvement in the preference estimator that could be achieved by optimistically selecting arm $i$ based on its reward estimation. In this way, it explicitly encourages global exploration of arms that reduce uncertainty in preference estimation while performing local exploration.

### v2

@@ 4 General Case with Hidden Preference
Figure 2: A scenario of user’s preference feedback is not explicitly provided (hidden preference).
@@
We first consider a more practical but more challenge scenario where only feedback on the reward and overall reward is observable, while preference feedback is hidden. For instance, in hotel surveys, customers often provide ratings on specific objectives (e.g., price, location, environment, amenities) along with an overall rating (as depicted in Fig. 2). In such cases, user preferences can be inferred from the latent relationship between the overall rating and the individual objective ratings. Formally, at each round $t$, the learner selects an arm $a_{t}\!\in\!\mathcal{A}_{t}^{n}$ for each user $n$, and observes the reward vector $\boldsymbol{r}_{a_{t}^{n}}\!\in\!\mathbb{R}^{D}$, and the overall-reward score $g_{a_{t}^{n},t}$.
@@
Within this framework, we adhere to the original Assumption 1 regarding rewards. It is worth noting that, in many real-world applications like hotel rating systems, the overall rating often shares the same scale as individual objective ratings. Therefore, we assume in this problem that the bound on the overall reward is identical to that of the individual rewards. This introduces one additional assumption and one revised assumption, as outlined below:
Assumption 4.
For $t\in[T]$, $n\in[N]$, $a_{t}^{n}\in[K]$, the overall-reward score satisfies $g_{a_{t}^{n},t}\in[0,1]$.
Assumption 5.
@@
For $t\in[T]$, $n\in[N]$, the stochastic preference is bounded and satisfies $\|\boldsymbol{c}_{t}^{n}\|_{1}\leq 1$. Without loss of generality, we assume $\boldsymbol{c}_{t}^{n}(d)$ is $R$-sub-Gaussian [^1] By Hoeffding’s lemma, for any $X\in[a,b]$ almost surely, $X$ is a $R$-sub-Gaussian random variable with $R$ at most (ba)/2(b-a)/2., $\forall d\in[D]$.
As discussed in Section 3, policy adapting to user preference is crucial. To enable effective learning and decision-making in this setting, we propose a unified framework that centers around two key components:
  • Preference Estimation: Inferring the user preference vector from the observed bandits feedback.
  • Preference-Aware Optimization: Selecting arms based on preference estimate to align decisions with user intent.
These two components serve as the algorithmic backbone of our approach and remain consistent across both hidden and revealed preference settings. However, in the hidden preference case, each component faces unique technical challenges.
@@ 4.1 Unique Challenges
Random Mapping from $\boldsymbol{r}_{t}$ to $g_{t}$.
In the hidden preference case, for each user $n\in[N]$, the observed overall rewards are generated through a random mapping of rewards. Specifically, $g_{a_{t}^{n},t}=(\boldsymbol{\overline{c}^{n}}+\boldsymbol{\zeta}_{t}^{n})^{\top}\boldsymbol{r}_{a_{t}^{n},t}=\boldsymbol{\overline{c}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}+{\boldsymbol{\zeta}_{t}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}$, where $\boldsymbol{\zeta}_{t}^{n}=\boldsymbol{c}_{t}^{n}-\boldsymbol{\overline{c}}^{n}\in\mathbb{R}^{D}$ is an independent random noise vector. This formulation implies that the overall residual noise term $\zeta_{g,t}^{n}={\boldsymbol{\zeta}_{t}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}$ is no longer independent of the input. Consequently, standard regression models become infeasible for preference estimation, as they rely on the assumption that the residual noise in the output is independent of the input.
Additionally, the magnitude of overall residual noise is a monotonically non-decreasing function w.r.t each reward objective, i.e., $\|\zeta_{g,t}(\boldsymbol{r}_{i})\|\leq\|\zeta_{g,t}(\boldsymbol{r}_{j})\|$ iff $\boldsymbol{r}_{i}\preceq\boldsymbol{r}_{j}$. This property implies that arms with smaller reward vectors, which are often suboptimal in terms of utility, may paradoxically offer more reliable information for preference estimation due to their lower sensitivity to noise. This issue would also have important implications for the optimization strategy, as we will discuss in the next challenge. Thus, a tailored latent preference estimator is essential to mitigate the expanding error w.r.t the reward and ensure effective preference learning.
Local Exploration vs Global Exploration.
Unlike traditional bandit algorithms that focus on a single goal (e.g., identifying the arm with the highest reward), the uncertainty in both preference and reward, combined with the need to infer latent preference, introduces a novel trade-off challenge: balancing global exploration for better preference estimation and local exploration of arm rewards:
@@
Note that these two learning objectives may conflict, as arms with high rewards might lack sufficient information for latent preference learning and may even worsen degrade estimation performance (as we discussed in the first challenge). This can also be verified by Fig. 3, where 80 samples of $\boldsymbol{r}_{t},g_{t}]$ are collected by repeatedly pulling an arm, and preference $\boldsymbol{\hat{c}}$ is estimated using linear regression. Here, $\boldsymbol{c}_{t}$ at each step follows a Gaussian distribution with a mean of $[0.5,0.5]$. The results demonstrate that samples from suboptimal Arm-1 (Fig. 3a) significantly outperform those from Pareto-optimal Arm-2 (Fig. 3b) in preference estimation. This necessitates an exploration policy that effectively addresses both global and local learning objectives.
Algorithm 1 PRUCB with Hidden Preference (PRUCB-HP)
Parameters: $\alpha$, $\lambda$, $\beta_{t}$, $\omega$.
Initialization: $\boldsymbol{\hat{r}}_{i,1}\!\leftarrow\![0]^{D},N_{i,1}\!\leftarrow\!0,\forall i\!\in\![K]$;For each user $n\in[N]$: $\boldsymbol{\hat{c}}_{1}^{n}\!\leftarrow\![1/D]^{D}$, $\boldsymbol{V}_{0}^{n}\!\leftarrow\!\lambda\boldsymbol{I}$.
for $t=1,\cdots,T$ do
  for user $n\in[N]$ do
[…]
   Pull arm $a_{t}\!\leftarrow\!\operatorname*{arg\,max}_{i\in\mathcal{A}_{t}^{n}}({\hat{\boldsymbol{c}}_{t}^{n}})^{\top}\hat{\boldsymbol{r}}_{i,t}\!+\!B_{i,t}^{n,r}\!+\!B_{i,t}^{n,c}$.
   Observe reward $\boldsymbol{r}_{a_{t}^{n},t}$ and overall-reward $g_{a_{t}^{n},t}$.
   Updating:
   $w_{t}^{n}=\frac{\omega}{\|\boldsymbol{r}_{a_{t}^{n},t}\|_{2}^{2}}$, $\boldsymbol{V}^{n}_{t}=\boldsymbol{V}^{n}_{t-1}+w_{t}^{n}\boldsymbol{r}_{a_{t}^{n},t}\boldsymbol{r}_{a_{t}^{n},t}^{\top}$.
   $\boldsymbol{\hat{c}}^{n}_{t+1}=({\boldsymbol{V}_{t}^{n}})^{-1}\sum_{\ell=1}^{\top}w_{\ell}^{n}g_{a_{\ell}^{n},\ell}\boldsymbol{r}_{a_{\ell}^{n},\ell}$.
  end for
  Updating:
  $N_{i,t+1}=N_{i,t}+\sum_{n\in[N]}\mathds{1}_{\{a_{t}^{n}=i\}}$.
  $\hat{\boldsymbol{r}}_{i,t+1}\!=\!\frac{\hat{\boldsymbol{r}}_{i,t}N_{i,t}+\sum_{n=1}^{N}\boldsymbol{r}_{a_{t}^{n},t}\cdot\mathds{1}_{\{a_{t}^{n}=i\}}}{N_{i,t+1}},\forall i\in[K]$.
end for
@@ 4.2 Our Algorithm
To this end, we propose a novel PRUCB-HP method (Algorithm 1) involving two key designs for both preference estimation and preference-aware optimization as follows.
Key design I: WLS-Preference Estimator.
For each user $n\in[N]$, as we have seen before, the randomness of preference $\boldsymbol{c}_{t}^{n}$ leads to the overall residual noise $\zeta_{g,t}^{n}$ be a function w.r.t, input reward $\boldsymbol{r}_{t}^{n}$. Moreover, larger input rewards $\boldsymbol{r}_{t}^{n}$ result in greater corruption from the residual noise. To resolve this, we employ a weighted least-squares (WLS) estimator for preference learning. Specifically, our algorithm assigns a weight $w_{t}^{n}$ to each observed sample and estimates the unknown preference using weighted ridge regression:
$\hat{\boldsymbol{c}}_{t}^{n}\xleftarrow{}\operatorname*{arg\,min}_{\boldsymbol{c}\in\mathbb{R}^{D}}\lambda\|\boldsymbol{c}\|_{2}^{2}+\sum_{\ell=1}^{t-1}w_{\ell}^{n}({\boldsymbol{c}}^{\top}\boldsymbol{r}_{a_{\ell}^{n},\ell}-g_{a_{\ell}^{n},\ell})^{2},$
@@
$\boldsymbol{\hat{c}}_{t}^{n}=({\boldsymbol{V}_{t-1}}^{n})^{-1}\sum_{\ell=1}^{t-1}w_{\ell}^{n}g_{a_{\ell}^{n},\ell}\boldsymbol{r}_{a_{\ell}^{n},\ell},$ (3)
where the Gram matrix $\boldsymbol{V}_{t-1}^{n}=\lambda\boldsymbol{I}+\sum_{\ell=1}^{t-1}w_{\ell}^{n}\boldsymbol{r}_{a_{\ell}^{n},\ell}\boldsymbol{r}_{a_{\ell}^{n},\ell}^{\top}$.
Inspired by Zhou, Gu, and Szepesvari (2021) using the inverse of the noise variance as weight for tight variance-dependent regret guarantee, we define the weight as the inverse of squared $\ell_{2}$-norm of the reward: $w_{t}^{n}=\omega/\|\boldsymbol{r}_{a_{t}^{n},t}\|_{2}^{2}$, where $\omega>0$ is a threshold parameter guaranteeing $w_{t}^{n}\geq 1$. Intuitively, it ensures samples with high rewards will be assigned smaller weights to reduce the influence of potentially large residual noises, while samples with low rewards receive larger weights to ensure their contribution to the estimation.
To see how our choice of weight can tackle the random mapping issue, we first define $\boldsymbol{r}_{a_{t}^{n},t}^{\prime}=\sqrt{w_{t}^{n}}\boldsymbol{r}_{a_{t}^{n},t}$, $g_{a_{t}^{n},t}^{\prime}=\sqrt{w_{t}^{n}}g_{a_{t}^{n},t}$, then the original formula Eq. 1 can be rewrite as
$\sqrt{w_{t}^{n}}\cdot g_{a_{t}^{n},t}=\sqrt{w_{t}^{n}}\cdot{\boldsymbol{c}_{t}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}=\sqrt{w_{t}^{n}}\cdot(\boldsymbol{\overline{c}}^{n}+\boldsymbol{\zeta_{t}}^{n})^{\top}\boldsymbol{r}_{a_{t}^{n},t}$
$\implies g_{a_{t}^{n},t}^{\prime}={\boldsymbol{\overline{c}}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}^{\prime}+\sqrt{w_{t}^{n}}\cdot{\boldsymbol{\zeta}_{t}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}.$ (4)
For term $\sqrt{w_{t}^{n}}\boldsymbol{\zeta}_{t}^{\top}\boldsymbol{r}_{a_{t}^{n},t}$, we have the following lemma:
@@
For any $n\in[N]$, the random variable $\sqrt{w_{t}^{n}}{\boldsymbol{\zeta}_{t}^{n}}^{\top}\boldsymbol{r}_{a_{t}^{n},t}$ is sub-Gaussian with constant $R^{\prime}=\sqrt{\omega}R$.
The proof is available in Appendix D.1. By above lemma, we observe that with the designed weight, the original random mapping regression problem is transferred into a new formula as (4). Specifically, the output $g_{a_{t}^{n},t}^{\prime}$ is mapped from $\boldsymbol{r}_{a_{t}^{n},t}^{\prime}$ via a fixed vector $\boldsymbol{\overline{c}}^{n}$ with a normed $R^{\prime}$-sub-Gaussian residual noise, where $R^{\prime}=\sqrt{\omega}R$, independent of the input.
Key design II: Dual-Exploration Policy.
As discussed earlier, there is a new global-local exploration dilemma in our setting. On the one hand, the algorithm must focus on local exploration by selecting optimistically profitable arms to discover better ones. Simultaneously, it must globally explore diverse arms to gather information about the relationship between $\boldsymbol{r}_{t}$ and $g_{t}$ for modeling the hidden preference.
To resolve this, we design an optimistic dual-exploration policy by incorporating a preference-driven bonus and reward-driven bonus under preference-aware optimization framework for trade-off. The optimistic policy is defined as
$a_{t}\leftarrow\operatorname*{arg\,max}_{i\in\mathcal{A}_{t}^{n}}({\hat{\boldsymbol{c}}_{t}^{n}})^{\top}\hat{\boldsymbol{r}}_{i,t}+B_{i,t}^{n,r}+B_{i,t}^{n,c},$ (5)
where $B_{i,t}^{n,r}$ and $B_{i,t}^{n,c}$ are the dual-exploration bonus terms for user $n$. We detail the design of these bonus terms below and will later theoretically demonstrate in Section 4.3 how they establish a tight UCB for the expected overall reward, ensuring the effectiveness of the optimistic policy in (5).
Reward Bonus $B_{i,t}^{n,r}$. The reward bonus term explicitly encourages local exploration of arms with potentially high rewards, for the principle of optimism in face of uncertainty. Specifically, the bonus $B_{i,t}^{n,r}$ is formulated as a reward uncertainty-aware regularization term:
$B_{i,t}^{n,r}=\rho_{i,t}^{\alpha}\|\hat{\boldsymbol{c}}_{t}^{n}\|_{1}.$ (6)
@@
Preference Bonus $B_{i,t}^{n,c}$. The preference bonus term aims to encourage the exploration of arms that reduce uncertainty in preference estimation. In previous bandit studies (Abbasi-Yadkori, Pál, and Szepesvári 2011; Zhao et al. 2020; He et al. 2022) involving linear coefficient ($\theta^{*}$) learning, it has been shown that $\beta\|\boldsymbol{x}_{i}\|_{\boldsymbol{V}^{-1}}$ provides a tight confidence bonus for the payoff of arm $i$, where $\beta$ is the confidence set radius for coefficient estimation, and $\boldsymbol{x}_{i}$ is the observable arm feature. In information theory, $\|\boldsymbol{x}_{i}\|_{\boldsymbol{V}^{-1}}$ also reflects entropy reduction in the model posterior, and is used to measure the estimator uncertainty improvement contributed by the chosen action $\boldsymbol{x_{i}}$ (Li et al. 2010).
However, such design is not feasible in our setting, as the exact reward $\boldsymbol{r}_{a_{t}^{n},t}$ is revealed only after pulling the arm $a_{t}^{n}$, making the actual information gain $\|\boldsymbol{r}_{a_{t}^{n},t}\|_{{(\boldsymbol{V}^{n}_{t-1})}^{-1}}$ from arm $a_{t}^{n}$ unpredictable beforehand. To resolve this problem, we introduce a pseudo information gain term, defined as $\|\boldsymbol{\hat{r}}_{i,t}+\rho_{i,t}^{\alpha}\boldsymbol{e}\|_{{(\boldsymbol{V}^{n}_{t-1})}^{-1}}$, where $\rho_{i,t}^{\alpha}$ is the standard Hoeffding bonus. And then the preference bonus is set as
$B_{i,t}^{n,r}=\beta_{t}\cdot\|\boldsymbol{\hat{r}}_{i,t}+\rho_{i,t}^{\alpha}\boldsymbol{e}\|_{{(\boldsymbol{V}^{n}_{t-1})}^{-1}},$ (7)
where $\beta_{t}$ is the confidence radius of the preference estimate we will give in Lemma 3. Intuitively, this pseudo information gain term captures the potential improvement of preference estimator could achieve by optimistically selecting arm $i$ based on its reward estimation. In this way, it explicitly encourages global exploration of arms that reduce uncertainty in preference estimation while performing local exploration.